OpenStudy (anonymous):

Use the discriminant to determine how many x-intercepts the graph of the equation has: y= -4x^2+3x-2 A.)None B.)One C.)Two D.)Three

OpenStudy (zzr0ck3r):

b^2-4ac =?

OpenStudy (anonymous):

I'm pretty sure that's the formula

OpenStudy (zzr0ck3r):

\(ax^2+bx+c\) then \(b^2-4ac = 0 \implies \text{one solution}\) \(b^2-4ac > 0 \implies \text{two solutions}\) \(b^2-4ac < 0 \implies \text{there are no real solutions}\) you need to figure out what \(b^2-4ac=\)

OpenStudy (zzr0ck3r):

\(a = -4\\ b = 3\\ c = -2\)

OpenStudy (zzr0ck3r):

what is \(3^2-4(-4)(-2)= \ ?\)

OpenStudy (zzr0ck3r):

ok good luck

OpenStudy (anonymous):

I don't get all this math crap. That's why I came here for answers.

OpenStudy (zzr0ck3r):

take it slow, and ask question if you dont understand, we prefer more questions than less/none.

OpenStudy (anonymous):

I just need to know how many intercepts the equation has. I have tried to figure it out and I can't. Math is my weakest subject.

OpenStudy (zzr0ck3r):

\(3^2-4(-4)(-2)= \ ?\)

OpenStudy (zzr0ck3r):

the answer to that will give the answer you seek...

OpenStudy (zzr0ck3r):

answer it and ill show you why....

OpenStudy (wolf1728):

If you encounter answers that are the square root of negative numbers, the equation WON'T intercept the x axis and will have NO real solutions.

OpenStudy (anonymous):

32

OpenStudy (anonymous):

@zzr0ck3r 32?

OpenStudy (zzr0ck3r):

3^2 -4(-4)(-2)= 9-32 < 0 so we have no real solutions because the discriminant is negative

OpenStudy (zzr0ck3r):

if it were positive we would have 2, if it were 0 we would have 1 3 will never be the answer...

OpenStudy (wolf1728):

Abbi - can you calculate this? b² - 4*a*c = 3² - (-4*-2)

OpenStudy (anonymous):

Thanks!